Weston Corporation just paid a dividend of $2.5 a share (i.e., D0 = $2.5). The dividend is expected to grow 12% a year for the next 3 years and then at 3% a year thereafter. What is the expected dividend per share for each of the next 5 years? Round your answers to two decimal places.

D1=?
D2=?
D3=?
D4=?
D5=?
Answer #1

Given D0 = $ 2.5 a share

It will grow at a rate of 12 % for next 3 years.

So,

D1 = 2.5 * 1.12 = $ 2.80 a share

D2 = 2.80 * 1.12 = $ 3.14 a share

D3 = 3.136 * 1.12 = $ 3.51 a share

After year 3, it is supposed to grow at a rate of 3%. So,

D4 = 3.51 * 1.03 = $ 3.62 a share

D5 = 3.62 * 1.03 = $ 3.73 a share